OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y=5x+10;(2,14) a. y=1/5x+4 b. y=-1/5x-4 c. y=5x-68 d. y=5x+4

OpenStudy (confxsedaf):

d.y=5x+4 you know that 5x+10 is y-intercept and slope you would do 10-14 to get m which is 4

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

@bmz yo niqqa this niqqa ugly af

OpenStudy (anonymous):

Could you help with a few more?? :/

OpenStudy (bmz):

lmao

OpenStudy (confxsedaf):

sure

OpenStudy (anonymous):

Write y =-2/3x+7 in standard form using integers. a.2x+3y =21 b. 3x-2y=21 c.-2x-3y=7 d. -2x+3y=7

OpenStudy (confxsedaf):

you would change -2 into 2 and times 7 by 3 to get 21 so a.2x+3y=21

OpenStudy (anonymous):

Thank you!